# Computational Design and Fabrication of Modular Robots with Untethered Control

### v1

@@ Appendix A Supplementary Information
. Stage $\epsilon_{approx}$ $w_{approx}$ $w_{barrier}$ $w_{map}$ $w_{mesh}$ $w_{fab-map}$ $w_{fab-shrink}$ $w_{fab}$ $w_{boundary}$ change topology coarse 0.035 1.0 $1e^{-5}$ 1.0 1.0 0.0 0.0 0.0 1.0 Yes fab 0.030 1.0 $1e^{-5}$ 0.0 1.0 1.0 200.0 1.0 1.0 No
@@
Algorithm Details: Exploring the Design and Control of Terrain Navigating Robots
@@
Given the skeletal graph $\Gamma(V,F)$ of the robot, we aim to optimize the positions($V^{\star}$) and the control gait of the active edges ($C_{e}^{\star}=A^{\star}_{e}(\sin(w^{\star}_{e}t+\phi^{\star}_{e}))$). We design various reward functions to capture different target behaviors. In the walking and swimming tasks, the reward is based on the total distance the robot travels in a given direction. For the task of climbing an inclined plane, the reward is determined by the height gained by the robot. The optimization algorithm proceeds in two steps. In the first step, we co-optimize the nodal positions and control parameters using the differential evolution algorithm [58], implemented in SciPy with a population size of $5$. This gives us an initial estimate of the vertices ($V^{0}$) and control gaits ($C_{e}^{0}$) of the robot. We subsequently refine the positions and control parameters through an alternating iterative approach (five iterations). Initially, the nodal positions are optimized with a fixed control gait, followed by optimizing the control parameters while maintaining a fixed skeleton. Optimization is again performed using a differential evolution method but with a larger population size of $30$. The frog-bots optimized using this method are presented in Fig. 6.

### v2

@@ Acknowledgments
. Stage $\epsilon_{approx}$ $w_{map}$ $w_{bij}$ $w_{boundary}$ $w_{tri}$ $w_{approx}$ $w_{fabmap}$ $w_{shrink}$ change topology map 0.035 1.0 $1e^{-5}$ 1.0 1.0 1.0 0.0 0.0 Yes fab 0.030 0.0 $1e^{-5}$ 1.0 1.0 1.0 1.0 200.0 No
@@
Algorithm Details for exploring the Design and Control of Terrain Navigating Robots
@@
Given the skeletal graph $\Gamma(V,F)$ of the robot, we aim to optimize the positions($V^{\star}$) and the control gait of the active edges ($C_{e}^{\star}=A^{\star}_{e}(\sin(w^{\star}_{e}t+\phi^{\star}_{e}))$). We design various reward functions to capture different target behaviors. In the walking and swimming tasks, the reward is based on the total distance the robot travels in a given direction. For the task of climbing an inclined plane, the reward is determined by the height gained by the robot. The optimization algorithm proceeds in two steps. In the first step, we co-optimize the nodal positions and control parameters using the differential evolution algorithm  (?), implemented in SciPy with a population size of $5$. This gives us an initial estimate of the vertices ($V^{0}$) and control gaits ($C_{e}^{0}$) of the robot. We subsequently refine the positions and control parameters through an alternating iterative approach (five iterations). Initially, the nodal positions are optimized with a fixed control gait, followed by optimizing the control parameters while maintaining a fixed skeleton. Optimization is again performed using a differential evolution method but with a larger population size of $30$. The frog-bots optimized using this method are presented in Figure 6.